OpenStudy (anonymous):

10m^(12/5)..write in radical form...

OpenStudy (lgbasallote):

\[\large x^{a/b} = \sqrt[b]{x^a}\]

OpenStudy (lgbasallote):

remember that rule

OpenStudy (lgbasallote):

also..when there is an exponent usually the one on its left is the only one affected

OpenStudy (lgbasallote):

therefore only m is raised to 12/5 and NOT 10

OpenStudy (lgbasallote):

can you go on from those info?

OpenStudy (lgbasallote):

that's a fun way to state it...

OpenStudy (lgbasallote):

do you get both of us @Lovely95 ??

OpenStudy (lgbasallote):

lol i love it when my avatar dominates the thread

OpenStudy (anonymous):

Yes, I get it! I think, I got \[10\times \sqrt[5]{m ^{12}}\]

OpenStudy (lgbasallote):

yes..but wait you stil need to simplify it!!

OpenStudy (anonymous):

no the question doesnt require simplification

OpenStudy (lgbasallote):

oh..well then you have the answer in your hands, young padawan..it is no longer blowing in the wind

OpenStudy (anonymous):

ahahaha ok thanks :D

OpenStudy (lgbasallote):

my pleasure ^_^ i would say my famous quote...but it seems inappropriate with this avatar so i'll just say you're welcome :)

OpenStudy (anonymous):

hahahah, ur funny ;)